7,800g=_ kg?
_ liters =240kl ?

could you please show me how to convert these problems?

Answer by Earlsdon(6098)

Remember! Kilo = 1,000
So, 1 kilogram = 1,000 grams or 1 gm = 1/1,000 Kg
(7,800g )/(1,000g/Kg) = 7.8 Kg
Similarly for liters.
1 Kiloliter = 1,000 liters or 1 liter = 1/1,000 Kl.
240 Kl = 240 Kl (1,000 l/Kl) = 240,000 liters